Question 1 You will establish whether or not current global fossil fuel combustion Co2 emissions can offset by planting trees? a) If all cropland was covered with slow-growing trees, how much fossil fuel Co2 emissions can you offset? b) How much with fast-growing trees? c) What happens when you use fast-growing trees on all arable land? d) Is this realistic? e) If this were to be done, what other issues would you need to consider? Some key factoids: o The surface of the Earth is 5 x 14 m? o cropland is~11% of the land surface o all arable land is ~45% of the land surface o fast-growing trees will produce ~10 metric tons of dry matter/hectare-year (with optimal placement and areal density) o slow-growing trees will produce ~1 metric tons of dry matterlhectare-year (with optimal placement and areal density) o typical dry matter is 0% carbon by mass o there are 10,000 m per hectare o current fossil fuel emissions: ~10 x10 metric tons carbon/year
Answer #1

a) Total cropland = 11% of land surface

thus (11/100)*5*1014=5.5*1013m2=5.5*109 hectartoe

Now slow growing trees fix = ~1 metric of dry matter / hectare / year

and 50 % of it is carbon. Thus actual offset = ( 50/100) * 1= 0.5 metric ton of carbon / hectare / year.

Thus amount of offset if all croplands replaced with slow growing trees will be 5.5*109*0.5= ~2.75*109 metric ton of carbon/hectare/year

b) Simillarly b) can be solved. Answer is 2.75*1010 metric ton of carbon / hectare/ year

c) Simillarly c) will be ~ 1.125*1011 metric ton of carbon / hectare /year.

d) No, I dont think this is realistic. It is not practical to replace all croplands and arable land with fast growing trees. Even if we did replace them, its not going to have very profound impact because the amount of emission is still very very high compared to the offset that will be achieved.

e) If itthis were to be done, serious issues will have to be taken in consideration. for example

1. How to take care of food production for the huge population?

2. What will happen to the different existing plant and animals which together form the ecosystem?

3. How to take up such large scale plantation and replacement?

4. To convince the people about such a thing.